# A flexible class of latent variable models for the analysis of antibody response data

### v1

@@ Abstract
Existing approaches to modelling antibody concentration data are mostly based on finite mixture models that rely on the assumption that individuals can be divided into two distinct groups: seronegative and seropositive. Here, we challenge this dichotomous modelling assumption and propose a latent variable modelling framework in which the immune status of each individual is represented along a continuum of latent seroreactivity, ranging from minimal to strong immune activation. This formulation provides greater flexibility in capturing age-related changes in antibody distributions while preserving the full information content of quantitative measurements. We show that the proposed class of models can accommodate a great variety of model formulations, both mechanistic and regression-based, and also includes finite mixture models as a special case. We demonstrate the advantages of this approach using malaria serology data and its ability to develop joint analyses across all ages that account for changes in transmission patterns. We conclude by outlining extensions of the proposed modelling framework and its relevance to other omics applications.
Keywords: serological data; latent variable model; antibody dynamics; mixture models; malaria; age-dependent immunity.
@@ 1 Introduction
Recent methodological advances have sought to extend GMMs by incorporating covariates, such as age or other demographic factors, which influence antibody concentrations and seropositivity status. For example, hardelid2008 used maternal age and country-of-birth to model variation in rubella antibody levels among newborns using a finite-mixture regression framework. Age-dependent mixture models are especially common in infectious disease applications: delfava2016 embedded a three-component mixture within a catalytic model to estimate age-specific immunity to varicella-zoster virus, while cox2022 used penalised splines to let seropositivity probabilities vary smoothly with age in dengue serology data. Similarly, kyomuhangi2021 combined antibody acquisition and catalytic models to allow for temporally varying seroconversion and boosting rates. These approaches reflect cumulative exposure risk by allowing mixing proportions or transition probabilities to evolve with age or other covariates. In parallel, efforts have been made to relax the normality assumption within each component. A key extension replaces Gaussian components with skew-normal distributions (azzalini1985), enabling the model to capture asymmetries commonly observed in antibody concentration data. Skew-normal mixtures retain analytical tractability while offering greater flexibility in representing the shape of seronegative and seropositive subpopulations (cabral2012).
@@
An alternative line of work has sought to avoid explicit dichotomisation through antibody acquisition (AA) models, in which age-dependent boosting and waning processes generate smooth trajectories of mean antibody levels (yman2016). Although these models provide a useful mechanistic link between age, exposure, and antibody dynamics, they typically rely on a log-Gaussian assumption and do not adequately accommodate the skewness and multimodality that characterise many serological datasets. Semiparametric methods have also been proposed that use ensemble machine learning to estimate age-antibody curves nonparametrically (arnold2017), but these target only the mean antibody level, precluding inference on other parameters of epidemiological interest. Our framework unifies the strengths of these approaches by accommodating both mechanistic interpretability through AA-type formulations and flexible modeling of the full heterogeneous distribution of antibody levels across age through a latent variable formulation.
@@
The remainder of the paper is organised as follows. In Section 2 we introduce the proposed latent variable modelling framework and outline alternative formulations for the latent seroreactivity distribution. Section 3 describes how age-dependent structure can be incorporated through both mechanistic and data-driven specifications. In Section 4 we detail the estimation strategy and validation procedures. Section 5 presents applications to malaria serology data, illustrating the flexibility and interpretability of the approach. We conclude in Section 6 with a discussion of the methodological implications and potential extensions.
@@ 2 The proposed latent variable modelling framework for serological data analysis
Figure 1 illustrates the distribution of latent serological activation levels. We use $T_{i}\in[0,1]$ to denote the latent variable representing the immunological activation state of individual $i$, for $i=1,\ldots,n$. In other words, $T$ represents a continuous measure of immune activation underlying the antibody response to the antigen of interest. Values of $T$ close to 0 correspond to a low-activation state, characterised by background or minimal serological activity, whereas values close to $1$ indicate a high-activation state reflecting strong or boosted antibody responses.
@@ 2.2 Modelling the latent state using a single-density parametric model
Specifically, the parameters $\alpha>0$ and $\beta>0$ determine the shape of the distribution. Larger values of $\alpha$ increase concentration near $T=1$, whereas larger values of $\beta$ shift mass toward $T=0$. Figure 2 illustrates how different Beta priors for $T$ influence the marginal distribution of the outcome $Y_{i}$ through the latent distribution of $T$. Different combinations of $(\alpha,\beta)$ can be interpreted in terms of epidemiological scenarios, with some configurations being more plausible than others depending on the context.
@@ 2.3 Modelling the latent state using mixture distributions
$f_{T}(t)=\sum_{c=1}^{3}\pi_{c}f_{c}(t_{i};\bm{\theta}_{c}),\quad t\in(0,1),$ (6)
where each $f_{c}(\cdot;\bm{\theta}_{c})$ is a component density defined on $(0,1)$ with parameters $\bm{\theta}_{c}$, and $\pi_{c}$ is the mixing weight for component $c$, subject to $\sum_{c=1}^{3}\pi_{c}=1$ and $\pi_{c}\geq 0$.
@@ 3.1 Single Beta distribution with age-dependent parameters
Here, $\mu(a)\in(0,1)$ denotes the expected antibody level at age $a$, while the precision parameter $\phi>0$ determines how concentrated the distribution is around the mean. The corresponding shape parameters are, in this case, $\alpha(a)=\mu(a)\phi$ and $\beta(a)=[1-\mu(a)]\phi$. Hence, the standard GMM framework here is recovered when $\phi=0$.
@@
This formulation introduces one fewer parameter than (3.1) and enables a straightforward interpretation of how age affects the expected value of $T$. The alternative parameterisation based on age-varying shape parameters in (3.1) offers greater flexibility to capture changes in the skewness of the distribution of $T$ across age but at the cost of reduced model interpretability.
@@ 3.2 Finite mixture models with age dependent mixing probabilities
An alternative way to introduce age dependence into the latent distribution of $T$ is through the mixing probabilities of a two-component Beta mixture model. In this formulation, the two Beta components represent subpopulations that differ in their propensity to mount an antibody response. The first component corresponds predominantly to individuals with low or absent immune activation, whereas the second represents those with a stronger or more sustained antibody response. This interpretation parallels that of classical finite mixture models directly fitted to the observed antibody measurements $Y$, but here it is formulated at the level of the latent variable $T$ so as to provide a more continuous representation of immune activity.
To ensure clear separation between the two Beta components, we impose the constraints $\alpha_{1}<1$ and $\beta_{1}>1$ for the first component, and $\alpha_{2}>1$ and $\beta_{2}<1$ for the second. This yields densities that are respectively concentrated near $T=0$ and $T=1$, providing a biologically interpretable distinction between the two subpopulations.
Let $\pi(a)$ denote the probability that an individual of age $a$ belongs to the antibody-reactive component. The distribution of $T$ can then be expressed as
@@
which arises from a constant hazard $\lambda$ for the onset of measurable immune activity. However, in our context, $\lambda$ should not be viewed as a seroconversion rate. Instead, it determines the speed at which higher levels of seroreactivity accumulate in the population as age increases. In other words, $\lambda$ controls how quickly the upper part of the latent immune state distribution gains weight with age, reflecting gradual changes in seroreactivity rather than transitions between discrete serostatus categories. We would expect the value of $\lambda$ to correlate with classical seroconversion rates, since both quantities reflect underlying exposure processes. Nevertheless, the values of $\lambda$ obtained from our latent mixture formulation are likely to be larger, because individuals can reach high levels of seroreactivity without necessarily becoming seropositive according to a binary threshold.
@@ 4 Inference
To estimate the parameters of the latent mixture model we use two procedures. First, we employ standard maximum likelihood; second, when many distributions need to be fitted or when fast approximate estimates are sufficient, we use a histogram-based fitting approach. The data here consist of antibody concentrations ${\bf Y}=[y_{1},\ldots,y_{n}]$. Recall that, conditionally on the latent immune state $T_{i}$, the outcome is assumed Gaussian with mean and variance that interpolate between the low seroreactivity and high seroreactivity extremes as in (2). The marginal density of $Y$ is therefore obtained by integrating over the distribution of the latent variable $T$,
$f(y;\bm{\theta})\;=\;\int_{0}^{1}\phi\!\left(y;\;(1-t)\mu_{0}+t\mu_{1},\;(1-t)\sigma_{0}^{2}+t\sigma_{1}^{2}\right)\,g_{T}(t;\bm{\psi})\,dt,$
where $\phi(\cdot;\mu,\sigma^{2})$ denotes the Gaussian density, $\bm{\theta}=(\mu_{0},\mu_{1},\sigma_{0},\sigma_{1})$ is the vector of conditional parameters, and $\bm{\psi}$ collects the parameters governing the distribution of $T$ (e.g. Beta shape parameters $\alpha,\beta$). The exact log-likelihood is then
$\ell(\bm{\theta},\bm{\psi})\;=\;\sum_{i=1}^{n}\log f(y_{i};\bm{\theta},\bm{\psi}).$
@@
$p_{j}(\bm{\theta},\bm{\psi})\;\approx\;f(m_{j};\bm{\theta},\bm{\psi})\,\Delta_{j},$
@@
Estimation is performed by comparing the empirical histogram $\widehat{f}_{j}$ with the model-implied densities $f(m_{j};\bm{\theta},\bm{\psi})$. Specifically, we minimise the $L_{2}$ criterion
$Q(\bm{\theta},\bm{\psi})\;=\;\sum_{j=1}^{J}\Bigl\{\widehat{f}_{j}-f(m_{j};\bm{\theta},\bm{\psi})\Bigr\}^{2},$ (12)
@@
A logarithmic variant can also be defined based on $\widehat{f}_{j}\cdot log(f(m_{j};\bm{\theta},\bm{\psi}))$, which corresponds to a discretization of the Kullback–Leibler divergence. An advantage of these approaches is that, unlike the exact likelihood, which requires evaluating the marginal distribution $f(y;\bm{\theta},\bm{\psi})$ at every observation, the histogram method requires evaluation only at the $J$ bin midpoints. Since typically $J\ll n$, this can lead to a substantial reduction in the number of integral computations. This is convenient when searching for a good fit or when finding initial values for maximum likelihood.
@@
The theoretical justification for this approach rests on the fact that, as the bin widths $\Delta_{j}$ shrink, the histogram density $\widehat{f}_{j}$ converges to the empirical distribution of the data, while $f(m_{j};\bm{\theta},\bm{\psi})$ converges to the model density at those points. In this limiting regime, minimisation of (12) defines a minimum–distance estimator (wolfowitz1957; beran1977) that minimises the $L_{2}$ distance between the empirical and model densities. Under weak regularity conditions, the logarithmic version converges to the same function as the log-likelihood. Consequently, as the partition is refined and $\max_{j}\Delta_{j}\to 0$, minimisers of (12) converge to the maximum likelihood estimator associated with the full continuous–data likelihood. The $L_{2}$ criterion has the advantage of being more robust, which is convenient for mixture models as they typically have many local minima and exhibit irregular behaviour around the parameter edges.
@@
We utilize the $L_{2}$ method for estimation of each time bin in Figure 6 and Figure 4, while for the other models we use maximum likelihood after finding the initial estimates from the method described above. To quantify uncertainty in finite samples, we rely on a parametric bootstrap, where replicated datasets are generated under the fitted model and the full estimation procedure is repeated to approximate the sampling distribution of the estimators.

### v2

@@ Abstract
Existing approaches to modelling antibody concentration data are mostly based on finite mixture models that rely on the assumption that individuals can be divided into two distinct groups: seronegative and seropositive. Here, we challenge this dichotomous modelling assumption and propose a latent variable modelling framework in which the immune status of each individual is represented along a continuum of latent seroreactivity, ranging from minimal to strong immune activation. This formulation provides greater flexibility in capturing age-related changes in antibody distributions while preserving the full information content of quantitative measurements. We show that the proposed class of models can accommodate a large variety of model formulations, both mechanistic and regression-based, and also includes finite mixture models as a special case. We also propose a computationally efficient $L_{2}$-based estimator as an alternative to maximum likelihood estimation, which substantially reduces computational cost, and we establish its consistency. Through a case study on malaria serology, we demonstrate how the flexibility of the novel framework enables joint analyses across all ages while accounting for changes in transmission patterns. We conclude by outlining extensions of the proposed modelling framework and its relevance to other omics applications.
Keywords: age-dependency; antibody dynamics; immunology; latent variable models; mixture models; malaria; serology.
@@ 1 Introduction
Recent methodological advances have sought to extend GMMs by incorporating covariates, such as age or other demographic factors, which influence antibody concentrations and seropositivity status. For example, hardelid2008 used maternal age and country-of-birth to model variation in rubella antibody levels among newborns using a finite-mixture regression framework. Age-dependent mixture models are especially common in infectious disease applications: delfava2016 embedded a three-component mixture within a catalytic model to estimate age-specific immunity to varicella-zoster virus, while cox2022 used penalised splines to let seropositivity probabilities vary smoothly with age in dengue serology data. Similarly, kyomuhangi2021 combined antibody acquisition and catalytic models to allow for temporally varying seroconversion and boosting rates. These approaches reflect cumulative exposure risk by allowing mixing proportions or transition probabilities to evolve with age or other covariates. In parallel, efforts have been made to relax the normality assumption within each component. A key extension replaces Gaussian components with skew-normal distributions (azzalini1985), enabling the model to capture asymmetries commonly observed in antibody concentration data. Skew-normal mixtures retain analytical tractability while offering greater flexibility in representing the shape of seronegative and seropositive subpopulations (dias_domingues2024).
@@
An alternative line of work has sought to avoid explicit dichotomisation through antibody acquisition (AA) models, in which age-dependent boosting and waning processes generate smooth trajectories of mean antibody levels (yman2016). Although these models provide a useful mechanistic link between age, exposure, and antibody dynamics, they typically rely on a log-Gaussian assumption and do not adequately accommodate the skewness and multimodality that characterise many serological datasets. Semiparametric methods have also been proposed that use ensemble machine learning to estimate age-antibody curves nonparametrically (arnold2017), but these target only the mean antibody level, precluding inference on other parameters of epidemiological interest. The proposed framework in this study unifies the strengths of these approaches by accommodating both mechanistic interpretability through AA-type formulations and flexible modeling of the full heterogeneous distribution of antibody levels across age through a latent variable formulation.
A key practical challenge in implementing the proposed framework is the computational burden associated with maximum likelihood estimation, which requires repeated numerical integration. To address this, we develop a computationally efficient alternative based on histogram-based $L_{2}$ minimum distance estimation that substantially reduces the computational burden. Our work builds on the $L_{2}$ error theory developed by scott2001parametric, which establishes the criteria under which $L_{2}$ estimators converge to the true parameter values. We extend this framework to our latent variable modelling framework and establish consistency of the resulting estimator under appropriate regularity conditions.
@@
The remainder of the paper is organised as follows. Section 2 introduces the proposed latent variable modelling framework, contrasts it with alternative latent formulations, and presents both single-density and mixture specifications for the latent seroreactivity distribution. Section 3 describes how age-dependent structure can be incorporated through mechanistic and data-driven parameterisations of both single Beta distributions and finite mixtures. Section 4 details maximum likelihood and histogram-based $L_{2}$ estimation procedures, establishes the consistency of the $L_{2}$ estimator, and describes model validation approaches. Section 5 presents a comprehensive simulation study comparing the statistical and computational efficiency of $L_{2}$ and maximum likelihood estimation across epidemiologically relevant scenarios. Section 6 demonstrates the framework through applications to malaria serology data for AMA1 and MSP1 antigens, including age-stratified model comparisons. We conclude in Section 7 with methodological implications and extensions to multivariate settings, spatial-temporal modelling, and high-dimensional omics applications.
@@ 2 The proposed latent variable modelling framework for serological data analysis
Figure 1 illustrates the distribution of latent immune activation across the population. We denote by $T_{i}\in[0,1]$ the latent variable representing the immunological activation state of individual $i$, for $i=1,\ldots,n$. This continuous measure quantifies the underlying immune response to the antigen of interest, with values near 0 corresponding to low-activation states characterised by minimal serological activity, and values near 1 indicating high-activation states reflecting strong or boosted antibody responses.
@@ 2.2 Modelling the latent state using a single-density parametric model
Specifically, the parameters $\alpha>0$ and $\beta>0$ determine the shape of the distribution. Larger values of $\alpha$ increase concentration near $T=1$, whereas larger values of $\beta$ shift mass toward $T=0$. Figure 2 illustrates how different Beta distributions for $T$ influence the marginal distribution of the outcome $Y_{i}$ through the latent distribution of $T$. Different combinations of $(\alpha,\beta)$ can be interpreted in terms of epidemiological scenarios, with some configurations being more plausible than others depending on the context.
@@ 2.3 Modelling the latent state using mixture distributions
$f_{T}(t)=\sum_{c=1}^{3}\pi_{c}f_{c}(t_{i};\psi_{c}),\quad t\in(0,1),$ (6)
where each $f_{c}(\cdot;\psi_{c})$ is a component density defined on $(0,1)$ with parameters $\psi_{c}$, and $\pi_{c}$ is the mixing weight for component $c$, subject to $\sum_{c=1}^{3}\pi_{c}=1$ and $\pi_{c}\geq 0$.
@@ 3.1 Single Beta distribution with age-dependent parameters
Here, $\mu(a)\in(0,1)$ denotes the expected antibody level at age $a$, while the precision parameter $\phi>0$ determines how concentrated the distribution is around the mean. The corresponding shape parameters are, in this case, $\alpha(a)=\mu(a)\phi$ and $\beta(a)=[1-\mu(a)]\phi$. Hence, the standard GMM framework here is recovered when $\phi=0$, as $T$ reduces to a Bernoulli variable such that $P[T=1;a]=\mu(a)$.
@@
We will show the application of this modelling approach in the case study of Section 6.
@@ 3.2 Finite mixture models with age dependent mixing probabilities
An alternative way to introduce age dependence into the latent distribution of $T$ is through the mixing probabilities of a two-component Beta mixture model. In this formulation, the two Beta components represent subpopulations that differ in their propensity to mount an antibody response. The first component corresponds predominantly to individuals with low or absent immune activation, whereas the second represents those with a stronger or more sustained antibody response. This interpretation parallels that of classical finite mixture models directly fitted to the observed antibody measurements $Y$, but here it is formulated at the level of $T$ representing the latent seroreactivity level, hence the model interpretation differs.
To avoid label switching between the two Beta components, we impose the constraints $\alpha_{1}<1$ and $\beta_{1}>1$ for the first component, and $\alpha_{2}>1$ and $\beta_{2}<1$ for the second. This yields densities that are respectively concentrated near $T=0$ and $T=1$, providing a biologically interpretable distinction between the two subpopulations.
Let $\pi(a)$ denote the probability that an individual of age $a$ belongs to the second component related to higher levels of seroreactivity. The distribution of $T$ can then be expressed as
@@
which arises from a constant hazard $\lambda$ for the onset of measurable immune activity. However, in our context, $\lambda$ should not be viewed as a seroconversion rate. Instead, it determines the speed at which higher levels of seroreactivity accumulate in the population as age increases. In other words, $\lambda$ controls how quickly the upper part of the latent immune state distribution gains weight with age, reflecting gradual changes in seroreactivity rather than transitions between discrete serostatus categories. We would expect the value of $\lambda$ to correlate with classical seroconversion rates, since both quantities reflect underlying exposure processes. Nevertheless, the values of $\lambda$ obtained from our latent mixture formulation are likely to be larger, because individuals can reach high levels of seroreactivity without necessarily becoming seropositive according to a predefined threshold in antibody levels.
@@ 4 Inference
Parameter estimation for the latent variable model requires integrating over the continuous distribution of the unobserved immune state $T$, which presents computational challenges for likelihood-based inference. We address this by considering and comparing two approaches: standard maximum likelihood estimation (MLE), and a computationally efficient histogram-based $L_{2}$ distance minimization method. This can be thought of a discretizations of the $L_{2}E$ method proposed by scott2001parametric. Although our main motivation for using the method is computational gain, the limiting case i.e the $L_{2}E$ method is also used for it’s robusntess scott2001parametric.
The observed data consist of antibody concentrations ${\bf Y}=[y_{1},\ldots,y_{n}]$ measured across individuals. Under the conditional specification in (2), the marginal density of $Y$ is obtained by integrating over the latent variable distribution:
$f(y;\vartheta)\;=\;\int_{0}^{1}\phi\!\left(y;\;(1-t)\mu_{0}+t\mu_{1},\;(1-t)\sigma_{0}^{2}+t\sigma_{1}^{2}\right)\,g_{T}(t;\psi)\,dt,$
where $\phi(\cdot;\mu,\sigma^{2})$ denotes the Gaussian density, $\vartheta=(\mu_{0},\mu_{1},\sigma_{0},\sigma_{1},\psi)$ encompasses all model parameters, and $\psi$ denotes the parameters governing the distribution of $T$ (e.g., Beta shape parameters $\alpha,\beta$). The exact log-likelihood is then
$\ell(\vartheta)\;=\;\sum_{i=1}^{n}\log f(y_{i};\vartheta).$
@@
$p_{j}(\vartheta)\;\approx\;f(m_{j};\vartheta)\,\Delta_{j},$ (12)
@@
Estimation is performed by comparing the empirical histogram $\widehat{f}_{j}$ with the model-implied densities $f(m_{j};\vartheta,\psi)$. Specifically, we minimise the $L_{2}$ criterion
$Q(\vartheta)\;=\;\sum_{j=1}^{J}\Bigl\{\widehat{f}_{j}-f(m_{j};\vartheta)\Bigr\}^{2},$ (13)
@@
A logarithmic variant can also be defined based on $\widehat{f}_{j}\cdot log(f(m_{j};\vartheta,\psi))$, which corresponds to a discretization of the Kullback–Leibler divergence. An advantage of these approaches is that, unlike the exact likelihood, which requires evaluating the marginal distribution $f(y;\vartheta,\psi)$ at every observation, the histogram method requires evaluation only at the $J$ bin midpoints. Since typically $J\ll n$, this can lead to a substantial reduction in the number of integral computations. This is especially convenient when assessing the goodness of fit of several different models and also finding starting values for MLE.
@@
The theoretical justification for this approach rests on the fact that, as the bin widths $\Delta_{j}$ shrink, the histogram density $\widehat{f}_{j}$ converges to the empirical distribution of the data, while $f(m_{j};\vartheta,\psi)$ converges to the model density at those points. In this limiting regime, minimisation of (13) defines a minimum–distance estimator (wolfowitz1957; beran1977) that minimises the $L_{2}$ distance between the empirical and model densities, i.e. the $L_{2}E$ method.
@@
We utilize the $L_{2}$ method for finding the initial estimates and input this as starting value for MLE in the case studies of Section 6. To quantify uncertainty in finite samples, we rely on a parametric bootstrap, where replicated datasets are generated under the fitted model and the full estimation procedure is repeated to approximate the sampling distribution of the estimators.
4.1 Consistency of the histogram-based $L_{2}$ estimator
We establish the consistency of our estimator using the asymptotic framework of $M$-estimation (vandervaartwellner1996). The estimator is defined by minimizing the integrated squared error (ISE) between the histogram and the model, a method rooted in the $L_{2}E$ theory proposed by scott2001parametric.
To align with the standard maximization notation of vandervaartwellner1996, we define the estimator $\widehat{\vartheta}_{n}$ as the maximizer of the negative squared distance. Let $\mathbb{M}_{n}(\vartheta)$ denote the empirical criterion and $\mathbb{M}(\vartheta)$ the population criterion:
$\mathbb{M}_{n}(\vartheta)\;$ $:=\;-\int_{I}\Bigl\{\widehat{f}_{n}(y)-f_{n}(y;\vartheta)\Bigr\}^{2}\,dy,$
$\mathbb{M}(\vartheta)\;$ $:=\;-\int_{I}\Bigl\{f_{0}(y)-f(y;\vartheta)\Bigr\}^{2}\,dy.$
Here, $\widehat{f}_{n}(y)$ is the histogram density estimator and $f_{n}(y;\vartheta)$ is the discretized (binned) model density in (12). The population criterion $\mathbb{M}(\vartheta)$ corresponds to the negative squared $L_{2}$ distance between the true density $f_{0}$ and the candidate model density $f(\cdot;\vartheta)$. As discussed by scott2001parametric, maximizing $\mathbb{M}(\vartheta)$ is equivalent to minimizing the $L_{2}$ distance to the true data-generating distribution, which provides inherent robustness properties compared to maximum likelihood.
In our practical implementation, the integral is computed over a finite range covered by bins. For the theoretical proof, we treat the integral over the full support $I$ (which may be finite or infinite).
We first state the needed assumptions.
Assumption 1.
Let $\vartheta_{0}\in\vartheta$ denote the true parameter such that $f_{0}(\cdot)=f(\cdot;\vartheta_{0})$. We posit the following regularity conditions:
  1. A1.
    Compactness. The parameter space $\vartheta$ is a compact subset of $\mathbb{R}^{p}$.
  2. A2.
    Identification. The true parameter $\vartheta_{0}$ is the unique maximizer of the population criterion $\mathbb{M}$ on $\vartheta$. Furthermore, for every open neighborhood $G$ of $\vartheta_{0}$, the maximum is well-separated:
    $\mathbb{M}(\vartheta_{0})\;>\;\sup_{\vartheta\in\vartheta\setminus G}\mathbb{M}(\vartheta).$
  3. A3.
    Regularity. The model family satisfies $\sup_{\vartheta\in\vartheta}\|f(\cdot;\vartheta)\|_{2}<\infty$. Additionally, the map $(y,\vartheta)\mapsto f(y;\vartheta)$ is measurable, and the true density $f_{0}\in L_{2}$.
  4. A4.
    Histogram Regime. The histogram partitions $\mathbb{R}$ into bins of equal width $h_{n}$. As $n\to\infty$, the bin width satisfies:
    $h_{n}\to 0\quad\text{and}\quad nh_{n}\to\infty.$
  5. A5.
    Smoothness. For every $\vartheta\in\vartheta$, the density $y\mapsto f(y;\vartheta)$ is absolutely continuous with first derivative $f^{\prime}(y;\vartheta)$. We assume the derivative is square-integrable uniformly over the parameter space:
    $\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}\;<\;\infty.$
Under these assumptions we can establish establish the consistency of the proposed $L_{2}$ estimator.
Theorem 4.1.
Under Assumptions A1.A5., let $\widehat{\vartheta}_{n}$ be the maximizer of $\mathbb{M}_{n}$ then $\widehat{\vartheta}_{n}\xrightarrow{p}\vartheta_{0}$ as $n\to\infty$.
Proof of the above theorem is given in Section A of the Appendix. In Section B of the Appendix we prove that the set of Assumptions A1.A5. hold for the proposed latent variable model in (2) under mild conditions on the mixing distribution $g_{T}$.